# DRoP: Distributionally Robust Pruning

### v2

@@ 1 Introduction
A recent study by (Pote et al., 2023) hints that data pruning may mitigate distributional bias in trained models, which is a well-established issue in AI systems concerning the performance disparity across classes (Dwork et al., 2012; Hardt et al., 2016b) or protected, minority groups of the population (e.g., race or gender). In Section 2, we conduct the first systematic evaluation of various pruning algorithms with respect to classification bias on a variety of benchmarks and conclude otherwise. For example, we find that Dynamic Uncertainty He et al. (2023) achieves superior average test performance with VGG-19 on CIFAR-100 and ResNet-50 on ImageNet but fails miserably in terms of worst-class accuracy. Thus, we argue that it is imperative to benchmark pruning algorithms using a more comprehensive suite of metrics that reflect classification bias, and to develop solutions that address distributional robustness directly.
@@
2 Data Pruning & Distributional Robustness
@@
Figure 2: Pruning Exacerbates Bias: Dynamic Uncertainty applied to CIFAR-100. See Appendix E for similar plots for other pruning methods and models. Left: Sorted class densities at different dataset density levels. We also report the minimum number of samples per class (SPC) at $10\%$ dataset density. Right: Full dataset test class-wise accuracy against dataset density. We also report the correlation coefficient between these two quantities across classes, averaged over $5$ dataset densities.
Data Pruning Exacerbates Bias.
Overall, all studied algorithms exhibit poor robustness to bias, although several of them improve ever so slightly over the full dataset. In particular, EL2N and GraNd achieve a relatively low classification bias, closely followed by Forgetting. At the same time, Forgetting has a substantially stronger average test accuracy compared to these three methods, falling short of Random only after pruning more than $60\%$ from CIFAR-100. As seen from additional plots in Appendix E, Forgetting produces more balanced datasets than EL2N and GraNd at low densities (Figure 14), and tends to prune “easier” classes more aggressively compared to all other methods (Figure 13). These two properties seem to be beneficial, especially when the available training data is scarce.
@@ 4 DRoP: Distributionally Robust Pruning
We are ready to propose our “robustness-aware” data pruning method, which consists in random subsampling according to carefully selected target class-wise sizes, incorporating lessons learned from our theoretical analysis. We propose to select the pruning fraction of each class based on its validation performance given a preliminary model $\psi$ trained on the whole dataset. Such a query model is still required by all existing data pruning algorithms to compute scores, so we introduce no additional resource overhead (see Appendix F for more discussion).
@@ Results.
While we scoped our study around robustness to classification bias, a vast majority of research is concerned with strong model performance across sensitive or minority groups within the distribution. One common benchmark dataset in this area is Waterbirds, which is a binary image classification of bird types where ground truth is spuriously correlated with image background (forest/water) for the majority of samples Sagawa* et al. (2020). Thus, standard optimization techniques produce models that often make predictions based on the background type rather than bird features, yielding poor accuracy for images from waterbird+forest or landbird+water groups. We use the original learning setup introduced by Sagawa* et al. (2020) to evaluate DRoP against other pruning baselines as well as CDB-W. Figure 9 shows that Random+DRoP and EL2N+DRoP achieve a significant improvement of the worst-group accuracy and other metrics. Therefore, as expected, our algorithm applies not only to reduce classification bias but also to improve group-wise robustness.
Figure 9: The average test performance of data pruning protocols against data density and measures of group-wise robustness (ResNet-50 on Waterbirds). Results averaged over 3 random seeds. Error bands represent min/max.
@@ Appendix A Implementation Details
Our empirical work encompasses three standard computer vision benchmarks (Table 1). All code is implemented in PyTorch (Paszke et al., 2017) and run on an internal cluster equipped with NVIDIA RTX8000 GPUs. The total runtime of the empirical work presented in this paper is approximately $45,000$ GPU hours.
@@
As shown in Table 1, we use the following model-dataset pairs: VGG-16 and VGG-19 (Simonyan & Zisserman, 2015) on CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), respectively, ResNet-18 (He et al., 2016) on TinyImageNet (MIT License) (Le & Yang, 2015), ImageNet pre-trained ResNet-50 on Waterbirds (Sagawa* et al., 2020) (MIT License), and ResNet-50 on ImageNet (Deng et al., 2009).
@@ Appendix D DRoP: Full Results
Appendix E Data Pruning: Class-Wise Analysis

### v4

# DRoP: Distributionally Robust Data Pruning

@@ 1 Introduction
A recent study by (Pote et al., 2023) hints that data pruning may mitigate distributional bias in trained models, which is a well-established issue in AI systems concerning the performance disparity across classes (Dwork et al., 2012; Hardt et al., 2016b) or protected, minority groups of the population (e.g., race or gender). In Section 3, we conduct the first systematic evaluation of various pruning algorithms with respect to classification bias—a phenomenon characterized by highly disparate model performance across different classes—on a variety of benchmarks and conclude otherwise. For example, we find that Dynamic Uncertainty He et al. (2023) achieves superior average test performance with VGG-19 on CIFAR-100 and ResNet-50 on ImageNet but fails miserably in terms of worst-class accuracy. Thus, we argue that it is imperative to benchmark pruning algorithms using a more comprehensive suite of metrics that reflect classification bias, and to develop solutions that address distributional robustness directly.
@@
2 Related Work
@@
3 Data Pruning is not Robust
Figure 1: Pruning Exacerbates Bias: Dynamic Uncertainty applied to CIFAR-100. See Appendix G for similar plots for other pruning methods and models. Left: Sorted class densities at different dataset density levels. We also report the minimum number of samples per class (SPC) at $10\%$ dataset density. Right: Full dataset test class-wise accuracy against dataset density. We also report the correlation coefficient between these two quantities across classes, averaged over $5$ dataset densities.
@@ 3 Data Pruning is not Robust
Overall, all studied algorithms exhibit poor robustness to bias, although several of them improve ever so slightly over the full dataset. In particular, EL2N and GraNd achieve a relatively low classification bias, closely followed by Forgetting. At the same time, Forgetting has a substantially stronger average test accuracy compared to these three methods, falling short of Random only after pruning more than $60\%$ from CIFAR-100. As seen from additional plots in Appendix G, Forgetting produces more balanced datasets than EL2N and GraNd at low densities (Figure 17), and tends to prune “easier” classes more aggressively compared to all other methods (Figure 16). These two properties seem to be beneficial, especially when the available training data is scarce.
@@ 5 DRoP: Distributionally Robust Pruning
We are ready to propose our “robustness-aware” data pruning method, which consists in random subsampling according to carefully selected target class-wise sizes, incorporating lessons learned from our theoretical analysis. We propose to select the pruning fraction of each class based on its validation performance given a preliminary model $\psi$ trained on the whole dataset. Such a query model is still required by all existing data pruning algorithms to compute scores, so we introduce no additional resource overhead (see Appendix E for more discussion).
@@ Results.
While we scoped our study around robustness to classification bias, a vast majority of research is concerned with strong model performance across sensitive or minority groups within the distribution. One common benchmark dataset in this area is Waterbirds, which is a binary image classification of bird types where ground truth is spuriously correlated with image background (forest/water) for the majority of samples Sagawa* et al. (2020). Thus, standard optimization techniques produce models that often make predictions based on the background type rather than bird features, yielding poor accuracy for images from waterbird+forest or landbird+water groups. We use the original learning setup introduced by Sagawa* et al. (2020) to evaluate DRoP against other pruning baselines as well as CDB-W. Figure 9 shows that Random+DRoP and EL2N+DRoP achieve a significant improvement of the worst-group accuracy and other metrics. For worst-class accuracy, we adopt a handful of prior art including Just Train Twice (Liu et al., 2021), Learning from Failure (Nam et al., 2020), Learning to Split (Bao & Barzilay, 2022), and Bias Amplification (Li et al., 2023); the performance of these algorithms in this setup is reported by Pezeshki et al. (2024), which we copied verbatim. As evident from Figure 9, Random+DroP is well on par with these more sophisticated techniques that operate on full datasets. Therefore, as expected, our algorithm applies not only to reduce classification bias but also to improve group-wise robustness.
Figure 9: The average test performance of data pruning protocols and existing baselines against measures of group-wise robustness (ResNet-50 on Waterbirds). The results of data pruning and CDB-W averaged over 3 random seeds. Error bands represent min/max. To conform with Sagawa* et al. (2020), for this dataset, we compute average accuracy as a sum of group accuracies weighted by the original training group proportions. This explains the sharp degradation of the average performance of DRoP-backed pruning at low densities ($d\leq 0.4$): these datasets are skewed towards minority groups that weigh much less than severely pruned majority groups. Please see Appendix F for an extended discussion.
@@ 6 Discussion
Acknowledgements
AV and JK acknowledge support through the NSF under award 1922658. Part of this work was done while JK was visiting the Centre Sciences de Donnees (CSD) at the Ecole Normale Superieure in Paris, France, and JK wishes to thank the CSD for their hospitality. This work was supported through the NYU IT High Performance Computing (HPC) resources, services, and staff expertise.
@@ Appendix A Implementation Details
Our empirical work encompasses three standard computer vision benchmarks (Table 1). All code is implemented in PyTorch (Paszke et al., 2017) and run on an internal cluster equipped with NVIDIA RTX8000 GPUs. The total runtime of the empirical work presented in this paper is approximately $45,000$ GPU hours. We make our code available at https://github.com/avysogorets/drop-data-pruning.
@@
As shown in Table 1, we use the following model-dataset pairs: VGG-16 and VGG-19 (Simonyan & Zisserman, 2015) on CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), respectively, ResNet-18 (He et al., 2016) on TinyImageNet (MIT License) (Le & Yang, 2015), ImageNet pre-trained ResNet-50 on Waterbirds (Sagawa* et al., 2020) (MIT License), and ResNet-50 on ImageNet (Deng et al., 2009). We also use a slight adaptation of Wide-ResNet-101 (Zagoruyko, 2016) on CIFAR-100 with a few downsampling layers removed.
@@ Appendix D DRoP: Full Results
[…]
Method DRoP Acc. CIFAR-10 CIFAR-100 TinyImageNet
$0.3$ $0.5$ $0.8$ $0.3$ $0.5$ $0.8$ $0.3$ $0.5$ $0.8$
Random avg $86.8$ $89.6$ $91.7$ $59.1$ $64.8$ $68.2$ $47.8$ $53.6$ $59.7$
worst $75.5$ $80.1$ $82.2$ $24.7$ $30.1$ $33.0$ $8.7$ $10.0$ $17.3$
avg $86.5$ $89.5$ $91.8$ $57.2$ $63.7$ $68.0$ $45.8$ $51.3$ $58.1$
worst $\mathbf{82.1}$ $\mathbf{85.9}$ $84.9$ $\mathbf{34.3}$ $\mathbf{45.3}$ $\mathbf{40.3}$ $\mathbf{16.2}$ $\mathbf{20.0}$ $16.9$
GraNd avg $87.6$ $91.9$ $92.7$ $40.4$ $58.5$ $69.3$ $30.4$ $48.7$ $60.0$
worst $77.0$ $83.6$ $85.1$ $12.3$ $31.7$ $39.0$ $1.0$ $12.4$ $15.8$
avg $85.9$ $91.6$ $92.6$ $35.5$ $55.9$ $68.7$ $29.9$ $48.6$ $59.4$
worst $78.0$ $83.8$ $84.7$ $15.0$ $30.1$ $36.3$ $4.7$ $10.7$ $16.7$
Forgetting avg $87.2$ $91.5$ $92.4$ $56.1$ $65.2$ $69.4$ $25.9$ $48.1$ $60.2$
worst $79.1$ $84.2$ $85.0$ $26.3$ $34.0$ $33.3$ $1.0$ $9.5$ $15.2$
avg $85.7$ $91.4$ $92.6$ $58.2$ $64.8$ $68.9$ $22.1$ $42.8$ $58.2$
worst $78.7$ $84.6$ $\mathbf{85.1}$ $32.0$ $34.0$ $38.3$ $1.3$ $12.7$ $\mathbf{18.0}$
Average avg $87.2$ $91.0$ $92.3$ $51.9$ $62.8$ $69.0$ $34.7$ $50.1$ $60.0$
worst $77.2$ $82.6$ $84.1$ $21.1$ $31.9$ $35.1$ $3.6$ $10.6$ $16.1$
avg $86.0$ $90.8$ $92.3$ $50.3$ $61.5$ $68.5$ $32.6$ $47.6$ $58.6$
worst $79.6$ $84.8$ $84.9$ $27.1$ $36.5$ $38.3$ $7.4$ $14.5$ $17.2$
Full Dataset avg $92.8$ $70.0$ $61.6$
worst $84.8$ $35.8$ $17.8$
Figure 13: The average test performance of various data pruning protocols against dataset density and measures of class robustness. Wide-ResNet-101 on CIFAR-100 with EL2N as Strategy. All results averaged over $3$ random seeds. Error bands represent min/max.
Larger Models.
Some previous studies argue that larger models enjoy better distributional robustness compared to smaller ones (Andreassen et al., 2022; Feuer et al., 2023). To see how model size could affect the applicability and effectiveness of DRoP, we train a Wide-ResNet-101 (Zagoruyko, 2016) with expansion factor of $2$ on CIFAR-100, which is an unusually high caliber architecture for such a relatively small dataset. In fact, we had to remove a few early downsampling layers from the original implementation of this model to account for the resolution of CIFAR-100. On full dataset, Wide-ResNet-101 with 125M parameters achieves a $4\%$ higher average and almost a $10\%$ higher worst-class accuracy compared to a VGG-19 with just 20M parameters (see Figures 11b and 13). Even in this more competitive setup, Random+DRoP continues to show considerable improvement in all robustness metrics over other pruning baselines as well as the full dataset (Figure 13).
Appendix E DRoP Ablation: Pre-Training Length
Appendix F ResNet-50 on Waterbirds
As stated in the caption of Figure 9, Sagawa* et al. (2020) compute average accuracy on Waterbirds as sum of group accuracy weighted by the group proportions in the training dataset, which is highly skewed towards majority classes (landbird+forest and waterbird+water, see Figure 15b). We adhere to this strategy in Figure 9 to ensure fair comparison with prior algorithms. Note, however, that this places DRoP at a disadvantage since it is designed to prune easier groups more aggressively and protect underperforming minority groups (see left barplots for each method in Figure 15b). Ultimately, this leads to a rapid degradation of the average accuracy at low dataset densities, which we observed in Figure 9 (left). In Figure 15a (bottom) we switch to a more standard accuracy computation as fraction of correctly classified samples regardless of group identities. In this framework, Random+DRoP remains strong relative to other methods. We should note that the test split is also slightly imbalanced with landbird+forest and landbird+water represented by $2255$ images each, and waterbird+forest and waterbird+water—by $642$.
Figure 15: (a): Average test accuracy against dataset density across different pruning methods. Top: Average accuracy computed as a weighted sum of group-wise accuracies according to training proportions. Bottom: Average accuracy is computed as a fraction of correctly classified samples of the test set. (b): Group sizes and group accuracy for all four groups in Waterbirds at $d=0.1$ for different pruning methods, as well as on the full dataset (bottom right barplots).
Appendix G Data Pruning: Class-Wise Analysis